OpenStudy (anonymous):

Which shows the fractions in order from least to greatest? five-twelfths, two-ninths, five-sixths

OpenStudy (texaschic101):

5/12 = 0.416 2/9 = 0.222 5/6 = 0.833 least to greatest : 2/9, 5/12, 5/6